# Huppert’s conjecture for finite simple exceptional groups of Lie type

### v1

@@ 6. Verifying Step 3: $G^{\prime}$ is isomorphic to $H$
By Step $2,$ we obtain that $G^{\prime}/M\cong H.$ Hence it suffices to show that $M=1.$ By way of contradiction, assume that $M$ is nontrivial. Let $K\leq M$ be a normal subgroup of $G^{\prime}$ such that $M/K$ is a chief factor of $G^{\prime}.$ Assume first that $M/K$ is nonabelian. It follows that $M/K\cong S^{k},$ where $k\geq 1$ and $S$ is a nonabelian finite simple group. By Lemma 2.1 and [2, Lemma 5], $M/K$ possesses a nontrivial irreducible character $\varphi\in{\mathrm{Irr}}(M/K)$ which extends to $\varphi_{0}\in{\mathrm{Irr}}(G^{\prime}).$ As $G^{\prime}/M\cong H,$ $G^{\prime}/M$ possesses an irreducible character $\chi$ of degree $|H|_{p},$ and hence by Gallagher’s Theorem, $\varphi_{0}\chi$ is an irreducible character of $G^{\prime}$ of degree $\varphi_{0}(1)|H|_{p}>|H|_{p}.$ As $G^{\prime}\unlhd G,$ we obtain that $\varphi_{0}(1)|H|_{p}$ must divide some degree of $G,$ which contradicts [25, Lemma 2.4]. Thus $M/K$ must be a minimal normal elementary abelian subgroup of $G^{\prime}/K.$ Let $C/K$ be the centralizer of $M/K$ in $G^{\prime}/K.$ As $M/K$ is a minimal normal elementary abelian subgroup of $G^{\prime}/K,$ we deduce that $M/K\leq C/K\unlhd G^{\prime}/K.$ In particular, $M\unlhd C\unlhd G^{\prime}.$ As $G^{\prime}/M$ is nonabelian simple, either $C=G^{\prime}$ or $C=M.$
@@
Assume $C=M.$ We have that $M/K$ is a self-centralizing minimal normal elementary abelian subgroup of $G^{\prime}/K$ and $G^{\prime}/M\cong H.$ Let $\chi\in{\mathrm{Irr}}(G^{\prime}/K)$ be such that $\chi(1)$ is the largest degree of $G^{\prime}/K.$ Applying Lemma 2.3 to $M/K\unlhd G^{\prime}/K,$ we have that $\chi(1)>b(H),$ which is a contradiction as $\chi(1)\in{\mathrm{cd}}(G^{\prime})$ must divide some degree of $H$ and thus $\chi(1)\leq b(H).$
@@
Assume $C=G^{\prime}$. Then $M/K=Z(G^{\prime}/K).$ Since $G^{\prime}$ is perfect, we obtain that $G^{\prime}/K$ is perfect and so $M/K=(G^{\prime}/K)^{\prime}\cap Z(G^{\prime}/K),$ which implies that $G^{\prime}/K$ is a central extension of $G^{\prime}/M$ (see [10, p. 629]). We now consider each case separately.
@@
Subcase $H\cong\textrm{F}_{4}(q),(q>2)$ or ${\rm E}_{8}(q)$. By [15, Theorem 5.1.4], the Schur multiplier of $G^{\prime}/M\cong H$ is trivial, so $G^{\prime}/K\cong G^{\prime}/M,$ which implies that $M/K$ is trivial, which is a contradiction.
@@
Subcase $H\cong{}^{2}\textrm{E}_{6}(q).$ By [15, Theorem 5.1.4], we know that the Schur multiplier of $G^{\prime}/M\cong H$ is cyclic of order $d=(3,q+1).$ If $d=1,$ then $G^{\prime}/K\cong H$ so $M/K$ is trivial, a contradiction. Hence $d=3.$ Assume first that $q>2.$ Then ${}^{2}{\rm E}_{6}(q)_{sc}$ is the universal covering group of $H$ so $G^{\prime}/K\cong{}^{2}{\rm E}_{6}(q)_{sc}.$ As $G^{\prime}\unlhd G$ and ${\mathrm{cd}}(G^{\prime}/K)\subseteq{\mathrm{cd}}(G^{\prime}),$ we deduce that every degree of $G^{\prime}/K$ must divide some degree of $G,$ which is impossible by Lemma 3.2$(x)$. Now assume that $q=2.$ We have that the Schur multiplier of ${}^{2}{\rm E}_{6}(2)$ is isomorphic to $\mathbb{Z}_{2}\times Z_{6}.$
If $G^{\prime}/K$ is isomorphic to $2\cdot{}^{2}{\rm E}_{6}(2)$ or $6\cdot{}^{2}{\rm E}_{6}(2),$ then $G^{\prime}/K$ possesses an irreducible character of degree $2^{7}\cdot 3^{5}\cdot 5^{2}\cdot 7^{2}\cdot 11\cdot 13\cdot 19$ which divides no degree of ${}^{2}{\rm E}_{6}(2)$ and if $G^{\prime}/K\cong 3\cdot{}^{2}{\rm E}_{6}(2),$ then $G^{\prime}/K$ has an irreducible character of degree $2^{11}\cdot 3^{3}\cdot 5\cdot 7^{2}\cdot 13\cdot 17\cdot 19$ Thus this case cannot happen.
The remaining cases can be argued similarly, we skip the details.
7. Verifying Step 4: $G=G^{\prime}\times C_{G}(G^{\prime}).$
We have shown that $G^{\prime}\cong H$. Let $C=C_{G}(G^{\prime}).$ Then $G/C$ is almost simple with socle $H.$ Assume that $G^{\prime}\times C<G.$ Then $G$ induces some outer automorphism on $G^{\prime}.$
@@
We first recall the set up in Section 3. Let $\mathcal{G}$ be a simple simply connected algebraic group in characteristic $p$ and $F$ is a Frobenius map such that $\mathcal{G}^{F}\cong H_{sc}=L,$ where $H\cong L/Z(L)$ is one of the simple groups in (1). Denote by $\mathcal{G}^{*}$ the dual group of $\mathcal{G}$ and let $F^{*}$ be the corresponding Frobenius automorphism. We have that $H^{*}:=(\mathcal{G}^{*})^{F^{*}}\cong H_{ad}.$
@@
Subcase $H\cong\textrm{F}_{4}(q),q>2.$ In this case, $H_{sc}\cong H\cong H_{ad}.$ Note that when $q$ is odd, $\textrm{Out}(H)$ is a cyclic group of order $f,$ consisting of field automorphisms, and when $p$ is even, $\textrm{Out}(H)$ is a cyclic group of order $2f$ with generator $\gamma$ such that $\gamma^{2}$ is a generator for the group of field automorphisms of $H.$
@@
Assume that $\sigma$ is a nontrivial field automorphism of $H.$ Let $\ell=\ell_{12f}(p)$ be a primitive prime divisor of $p^{12f}-1$ and let $s$ be a semisimple element of order $\ell$ in $(\mathcal{G}^{*})^{F^{*}}\cong H.$ We have that $|C_{(\mathcal{G}^{*})^{F^{*}}}(s)|=\Phi_{12}(q)$ and $H$ possesses an irreducible character $\chi_{s}$ with degree $|(\mathcal{G}^{*})^{F^{*}}:C_{(\mathcal{G}^{*})^{F^{*}}}(s)|_{p^{\prime}}=\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{8}.$ Let $\sigma^{*}$ be an automorphism of $\mathcal{G}^{*}$ induced from the field automorphism $\sigma.$ We obtain that $(\mathcal{G}^{*})^{\sigma^{*}}\cong{\rm F}_{4}(p^{b}),$ where $1\leq b<f,$ so that $\ell=\ell_{12f}(p)$ does not divide $|(\mathcal{G}^{*})^{\sigma^{*}}|,$ and hence by [6, Lemma 2.5(ii)], $\chi_{s}$ is not $\sigma$-invariant, and hence some proper multiple of $\chi_{s}(1)$ is a degree of $H,$ contradicting Lemma 3.1$(v)$.
@@
Now assume that $\sigma$ is not a field automorphism of $H.$ Then $p=2$ and $\sigma=\gamma^{k}$ for some odd integer $1\leq k\leq f,$ where $\gamma$ is a generator of $\mathop{\rm Out}\nolimits(H).$ Let $\tau=\sigma^{2}=\gamma^{2k}.$ If $\tau$ is nontrivial, then it is a field automorphism of $H,$ by applying the same argument as above, we obtain that $\chi_{s}$ is not $\tau$-invariant and so obviously $\chi_{s}$ is not $\sigma$-invariant, which leads to a contradiction. Hence $\sigma^{2}=1,$ which forces $f=k$ is odd. Thus $\sigma$ is a graph automorphism of order $2$ of ${\rm F}_{4}(2^{f}),$ where $f\geq 2$ is odd. By [17, Theorem 2.5(e)], $H$ has two unipotent characters labeled by the symbol $\phi_{2,16}^{\prime}$ and $\phi_{2,16}^{\prime\prime}$ with degree $\frac{1}{2}q^{13}\Phi_{4}\Phi_{8}\Phi_{12},$ which are fused under $\sigma.$ Therefore $G/C$ possesses an irreducible character of degree $q^{13}\Phi_{4}\Phi_{8}\Phi_{12}.$ However this contradicts Lemma 3.1$(vi)$ as $q$ is even. Therefore, $G=G^{\prime}C=G^{\prime}\times C,$ since $C\cap G^{\prime}=1,$ and hence $C\cong G/G^{\prime}$ is abelian.
@@
Subcase $H\cong{}^{2}\textrm{E}_{6}(q).$ Note that $\textrm{Out}(H)$ is a group of order $2df,$ where $d=(3,q+1).$ Assume first that $G/C$ possesses only inner and diagonal automorphism. It follows that $G/C\cong{}^{2}{\rm E}_{6}(q)_{ad},$ $3\mid(q+1)$ and $\sigma$ is a diagonal automorphism of $H.$ Thus we obtain that ${\mathrm{cd}}({}^{2}{\rm E}_{6}(q)_{ad})={\mathrm{cd}}(G/C)\subseteq{\mathrm{cd}}(G)={\mathrm{cd}}(H),$ which is impossible by Lemma 3.2$(x)$.
@@
Hence we can assume that $\sigma=\gamma\varrho,$ where $\varrho$ is a field automorphism and $\gamma$ is either trivial or a diagonal automorphism. Let $\ell=\ell_{18f}(p)$ be a primitive prime divisor of $p^{18f}-1$ and let $s$ be a semisimple element of order $\ell$ in $(\mathcal{G}^{*})^{F^{*}}.$ We have that $|C_{(\mathcal{G}^{*})^{F^{*}}}(s)|=\Phi_{18}$ and $\mathcal{G}^{F}$ possesses an irreducible character $\chi_{s}$ with degree $|(\mathcal{G}^{*})^{F^{*}}:C_{(\mathcal{G}^{*})^{F^{*}}}(s)|_{p^{\prime}}=|{}^{2}{\rm E}_{6}(q)_{ad}|_{p^{\prime}}/\Phi_{18}$ which acts trivially on its center, so $\chi_{s}\in{\mathrm{Irr}}(H).$ Let $\gamma^{*}$ be an automorphism of $\mathcal{G}^{*}$ induced from the field automorphism $\gamma.$ We obtain that $(\mathcal{G}^{*})^{\gamma^{*}}\cong{}^{2}{\rm E}_{6}(p^{b})_{ad},$ where $1\leq b<f,$ so $\ell=\ell_{18f}(p)$ does not divide $|(\mathcal{G}^{*})^{\gamma^{*}}|,$ and hence by [6, Lemma 2.5(ii)], we deduce that $\chi_{s}$ is not $\gamma$-invariant. As $\chi_{s}$ is invariant under the diagonal automorphisms, we deduce that $\chi_{s}$ is not $\sigma$-invariant and hence some proper multiple of $\chi_{s}(1)$ is a degree of $H,$ which is impossible. Thus, $G=G^{\prime}C=G^{\prime}\times C,$ since $C\cap G^{\prime}=1,$ and hence $C\cong G/G^{\prime}$ is abelian.
Subcase $H\cong\textrm{E}_{6}(q).$ Note that $\textrm{Out}(H)\cong\langle\gamma\rangle(\langle\varrho\rangle\times\langle\tau\rangle))$ is a group of order $2df,$ where $d=(3,q-1),$ $\gamma$ is a diagonal automorphism of order $d,$ $\varrho$ is a field automorphism of order $f$ and $\tau$ is a graph automorphism of order $2$ (see for example [7, Theorem 2.5.12]). Moreover $\tau$ inverts the cyclic group $\langle\gamma\rangle.$
Assume first that $G/C$ possesses only inner and diagonal automorphism. It follows that $G/C\cong{\rm E}_{6}(q)_{ad},$ $13\mid q-$ 1 and $\sigma$ is a diagonal automorphism of $H.$ Thus ${\mathrm{cd}}({\rm E}_{6}(q)_{ad})={\mathrm{cd}}(G/C)\subseteq{\mathrm{cd}}(G)={\mathrm{cd}}(H),$ which is impossible by Lemma 3.3$(x)$.
@@
Hence we assume next that $\sigma=\gamma\varrho,$ where $\varrho$ is a nontrivial field automorphism and $\gamma$ is either trivial or a diagonal automorphism. Let $\ell=\ell_{12f}(p)$ be a primitive prime divisor of $p^{12f}-1$ and let $s$ be a semisimple element of order $\ell$ in $K\cong{\rm F}_{4}(q)$ where $K\leq(H^{*})^{\prime}$ is chosen so that it is fixed under the involutory graph $\tau$ (see [7, Proposition 4.9.2]). We have that $|C_{(\mathcal{G}^{*})^{F^{*}}}(s)|=\Phi_{12}\Phi_{3}$ and $\mathcal{G}^{F}$ possesses an irreducible character $\chi_{s}$ with degree $|(\mathcal{G}^{*})^{F^{*}}:C_{(\mathcal{G}^{*})^{F^{*}}}(s)|_{p^{\prime}}=|{\rm E}_{6}(q)_{ad}|_{p^{\prime}}/\Phi_{12}\Phi_{3},$ which is trivial on the center as $s\in(H^{*})^{\prime}$ so $\chi_{s}\in{\mathrm{Irr}}(H).$ Let $\gamma^{*}$ be an automorphism of $\mathcal{G}^{*}$ induced from the field automorphism $\gamma.$ We obtain that $(\mathcal{G}^{*})^{\gamma^{*}}\cong{\rm E}_{6}(p^{b})_{ad},$ where $1\leq b<f,$ so $\ell=\ell_{12f}(p)$ does not divide $|(\mathcal{G}^{*})^{\gamma^{*}}|,$ and hence by [6, Lemma 2.5(ii)], we deduce that $\chi_{s}$ is not $\gamma$-invariant. As $\chi_{s}$ is invariant under the diagonal automorphisms, we deduce that $\chi_{s}$ is not $\sigma$-invariant and hence some proper multiple of $\chi_{s}(1)$ is a degree of $H,$ which is impossible.
@@
Now assume that $\sigma=\gamma\varrho\tau,$ where $\gamma$ is either trivial or a diagonal automorphism, $\varphi$ is a nontrivial field automorphism and $\tau$ is a graph automorphism. In this case, as the field and graph automorphisms commute by [7, Theorem 2.5.12], we can write $\sigma=\gamma\tau\varrho.$ Choose the semisimple element $s\in(H^{*})^{\prime}$ as in previous paragraph with $|s|=\ell_{12}(q).$ Since $s$ in $\tau$-invariant, [22, Corollary 2.5] yields that $\chi_{s}^{\tau}=\chi_{\tau(s)}=\chi_{s}.$ As $\chi_{s}$ is also $\gamma$-invariant by applying [22, Corollay 2.5] again, $\chi_{s}$ is $\gamma\tau$-invariant. Thus $\chi_{s}^{\sigma}=\chi_{s}^{\gamma\tau\varrho}=\chi_{s}^{\varrho}\neq\chi_{s}.$ Therefore $\chi_{s}$ is not $\sigma$-invariant, which leads to a contradiction.

### v3

@@ 6. Verifying Step 3: $G^{\prime}$ is isomorphic to $H$
We prove the following result in this section.
Proposition 6.1.
Let $G$ be a finite group and let $H$ be one of the finite simple groups in (1). Let $M\leq G^{\prime}$ be a normal subgroup of $G$ such that $G^{\prime}/M$ is a chief factor of $G$. Then $G^{\prime}\cong H.$
Proof.
@@
By Propositions 4.1 and 5.1, $G^{\prime}$ is perfect and $G^{\prime}/M\cong H.$ It suffices to show that $M=1.$ By way of contradiction, assume that $M$ is nontrivial. Let $K\leq M$ be a normal subgroup of $G^{\prime}$ such that $M/K$ is a chief factor of $G^{\prime}.$ Assume first that $M/K$ is nonabelian. Then $M/K\cong S^{k},$ where $k\geq 1$ and $S$ is a nonabelian finite simple group. By Lemmas 2.1 and 2.2, $M/K$ possesses a nontrivial irreducible character $\varphi\in{\mathrm{Irr}}(M/K)$ which extends to $\varphi_{0}\in{\mathrm{Irr}}(G^{\prime}).$ As $G^{\prime}/M\cong H,$ it possesses an irreducible character $\chi$ of degree $|H|_{p},$ and hence by Gallagher’s Theorem, $\varphi_{0}\chi$ is an irreducible character of $G^{\prime}$ of degree $\varphi_{0}(1)|H|_{p}>|H|_{p}.$ As $G^{\prime}\unlhd G,$ $\varphi_{0}(1)|H|_{p}$ divides some degree of $G,$ contradicting [29, Lemma 2.4]. Thus $M/K$ is a minimal normal elementary abelian subgroup of $G^{\prime}/K.$ Let $C/K$ be the centralizer of $M/K$ in $G^{\prime}/K.$ As $M/K$ is a minimal normal elementary abelian subgroup of $G^{\prime}/K,$ we deduce that $M/K\leq C/K\unlhd G^{\prime}/K.$ In particular, $M\unlhd C\unlhd G^{\prime}.$ As $G^{\prime}/M$ is nonabelian simple, either $C=G^{\prime}$ or $C=M.$
@@
Assume $C=M.$ Then $M/K$ is a self-centralizing minimal normal elementary abelian subgroup of $G^{\prime}/K$ and $G^{\prime}/M\cong H.$ Let $\chi\in{\mathrm{Irr}}(G^{\prime}/K)$ be such that $\chi(1)$ is the largest degree of $G^{\prime}/K.$ Applying Lemma 2.5 to $M/K\unlhd G^{\prime}/K,$ we obtain $\chi(1)>b(H).$ Now $\chi(1)\in{\mathrm{cd}}(G^{\prime})$ divides some degree of $H$ and thus $\chi(1)\leq b(H)$, a contradiction.
@@
Assume $C=G^{\prime}$. Then $M/K=Z(G^{\prime}/K).$ Since $G^{\prime}$ is perfect, $G^{\prime}/K$ is perfect and so $M/K=(G^{\prime}/K)^{\prime}\cap Z(G^{\prime}/K),$ which implies that $G^{\prime}/K$ is a central extension of $G^{\prime}/M$ (see [13, p. 629]). We now consider each case separately.
@@
Subcase $H\cong\textrm{F}_{4}(q),(q>2)$ or ${\rm E}_{8}(q)$. By [18, Theorem 5.1.4], the Schur multiplier of $G^{\prime}/M\cong H$ is trivial, so $G^{\prime}/K\cong G^{\prime}/M,$ which implies that $M/K$ is trivial, a contradiction.
@@
Subcase $H\cong{}^{2}\textrm{E}_{6}(q).$ By [18, Theorem 5.1.4], the Schur multiplier of $G^{\prime}/M\cong H$ is cyclic of order $d=(3,q+1).$ If $d=1,$ then $G^{\prime}/K\cong H$ so $M/K$ is trivial, a contradiction. Hence $d=3.$ Assume first that $q>2.$ Then ${}^{2}{\rm E}_{6}(q)_{sc}$ is the universal covering group of $H$ so $G^{\prime}/K\cong{}^{2}{\rm E}_{6}(q)_{sc}.$ As $G^{\prime}\unlhd G$ and ${\mathrm{cd}}(G^{\prime}/K)\subseteq{\mathrm{cd}}(G^{\prime}),$ every degree of $G^{\prime}/K$ divides some degree of $G,$ which is impossible by Lemma 3.1$(x)$. Now assume that $q=2.$ The Schur multiplier of ${}^{2}{\rm E}_{6}(2)$ is isomorphic to $\mathbb{Z}_{2}\times\mathbb{Z}_{6}.$ If $G^{\prime}/K$ is isomorphic to $2\cdot{}^{2}{\rm E}_{6}(2)$ or $6\cdot{}^{2}{\rm E}_{6}(2),$ then $G^{\prime}/K$ possesses an irreducible character of degree $2^{7}\cdot 3^{5}\cdot 5^{2}\cdot 7^{2}\cdot 11\cdot 13\cdot 19$ which divides no degree of ${}^{2}{\rm E}_{6}(2)$. If $G^{\prime}/K\cong 3\cdot{}^{2}{\rm E}_{6}(2),$ then $G^{\prime}/K$ has an irreducible character of degree $2^{11}\cdot 3^{3}\cdot 5\cdot 7^{2}\cdot 13\cdot 17\cdot 19$ Thus this case cannot occur.
The remaining cases can be argued similarly, we skip the details. ∎
7. Verifying Step 4: $G=G^{\prime}\times C_{G}(G^{\prime})$
@@
In this section, we complete the proof of the main theorem. We first recall the setup in Section 3. Let $\mathcal{G}$ be a simple simply connected algebraic group in characteristic $p$ and $F$ is a Steinberg endomorphism such that $\mathcal{G}^{F}\cong H_{sc}=L,$ where $H\cong L/Z(L)$ is one of the simple groups in (1). Denote by $\mathcal{G}^{*}$ the dual group of $\mathcal{G}$ and let $F^{*}$ be the corresponding Frobenius automorphism. Then $H^{*}:=(\mathcal{G}^{*})^{F^{*}}\cong H_{ad}.$
Proposition 7.1.
Let $G$ be a finite group and let $H$ be one of the finite simple groups in (1). Suppose that ${\mathrm{cd}}(G)={\mathrm{cd}}(H).$ Then $G\cong G^{\prime}\times C_{G}(G^{\prime}).$
Proof.
By Proposition 6.1, $G^{\prime}\cong H$. Let $C=C_{G}(G^{\prime}).$ Then $G/C$ is almost simple with socle $H$ and $G^{\prime}\cap C=1$. If $G=G^{\prime}C$, then $G\cong G^{\prime}\times C$ and we are done. So, assume by contradiction that $G^{\prime}\times C<G.$ Then $G$ induces some outer automorphism, say $\sigma$, on $H.$
@@
Subcase $H\cong\textrm{F}_{4}(q),q>2.$ In this case, $H_{sc}\cong H\cong H_{ad}.$ When $p$ is odd, $\textrm{Out}(H)$ is a cyclic group of order $f,$ consisting of field automorphisms, and when $p$ is even, $\textrm{Out}(H)$ is a cyclic group of order $2f$ with generator $\gamma$ such that $\gamma^{2}$ is a generator for the group of field automorphisms of $H.$
@@
Assume that $\sigma$ is a nontrivial field automorphism of $H.$ Let $\ell=\ell_{12f}(p)$ be a primitive prime divisor of $p^{12f}-1$ and let $s$ be a semisimple element of order $\ell$ in $(\mathcal{G}^{*})^{F^{*}}\cong H.$ Now $|C_{(\mathcal{G}^{*})^{F^{*}}}(s)|=\Phi_{12}(q)$ and $H$ possesses an irreducible character $\chi_{s}$ with degree
$|(\mathcal{G}^{*})^{F^{*}}:C_{(\mathcal{G}^{*})^{F^{*}}}(s)|_{p^{\prime}}=\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{8}.$
Let $\sigma^{*}$ be an automorphism of $\mathcal{G}^{*}$ induced from the field automorphism $\sigma.$ Then $(\mathcal{G}^{*})^{\sigma^{*}}\cong{\rm F}_{4}(p^{b}),$ where $1\leq b<f,$ so $\ell=\ell_{12f}(p)$ does not divide $|(\mathcal{G}^{*})^{\sigma^{*}}|$. By [8, Lemma 2.5(ii)], $\chi_{s}$ is not $\sigma$-invariant, and hence some proper multiple of $\chi_{s}(1)$ is a degree of $H,$ contradicting Lemma 3.2$(v)$.
@@
Now assume that $\sigma$ is not a field automorphism of $H.$ Then $p=2$ and $\sigma=\gamma^{k}$ for some odd integer $1\leq k\leq f,$ where $\gamma$ is a generator of $\mathop{\rm Out}\nolimits(H).$ Let $\tau=\sigma^{2}=\gamma^{2k}.$ If $\tau$ is nontrivial, then it is a field automorphism of $H$. By applying the same argument as above, we obtain that $\chi_{s}$ is not $\tau$-invariant and so obviously $\chi_{s}$ is not $\sigma$-invariant, a contradiction. Hence $\sigma^{2}=1,$ which forces $f=k$ is odd. Thus $\sigma$ is a graph automorphism of order $2$ of ${\rm F}_{4}(2^{f}),$ where $f\geq 2$ is odd. By [21, Theorem 2.5(e)], $H$ has two unipotent characters labeled by the symbol $\phi_{2,16}^{\prime}$ and $\phi_{2,16}^{\prime\prime}$ with degree $\frac{1}{2}q^{13}\Phi_{4}\Phi_{8}\Phi_{12},$ which are fused under $\sigma.$ Therefore $G/C$ possesses an irreducible character of degree $q^{13}\Phi_{4}\Phi_{8}\Phi_{12}.$ However this contradicts Lemma 3.2$(vi)$ as $q$ is even.
@@
Subcase $H\cong{}^{2}\textrm{E}_{6}(q).$ Note that $\textrm{Out}(H)$ is a group of order $2df,$ where $d=(3,q+1).$ Assume first that $G/C$ possesses only inner and diagonal automorphism. It follows that $G/C\cong{}^{2}{\rm E}_{6}(q)_{ad},$ $3\mid(q+1)$ and $\sigma$ is a diagonal automorphism of $H.$ Thus ${\mathrm{cd}}({}^{2}{\rm E}_{6}(q)_{ad})={\mathrm{cd}}(G/C)\subseteq{\mathrm{cd}}(G)={\mathrm{cd}}(H),$ which is impossible by Lemma 3.1$(x)$.
@@
Hence we can assume that $\sigma=\gamma\varrho,$ where $\varrho$ is a field automorphism and $\gamma$ is either trivial or a diagonal automorphism. Let $\ell=\ell_{18f}(p)$ be a primitive prime divisor of $p^{18f}-1$ and let $s$ be a semisimple element of order $\ell$ in $(\mathcal{G}^{*})^{F^{*}}.$ Now $|C_{(\mathcal{G}^{*})^{F^{*}}}(s)|=\Phi_{18}$ and $\mathcal{G}^{F}$ possesses an irreducible character $\chi_{s}$ with degree
$|(\mathcal{G}^{*})^{F^{*}}:C_{(\mathcal{G}^{*})^{F^{*}}}(s)|_{p^{\prime}}=|{}^{2}{\rm E}_{6}(q)_{ad}|_{p^{\prime}}/\Phi_{18}$
@@
which acts trivially on its center, so $\chi_{s}\in{\mathrm{Irr}}(H).$ Let $\varrho^{*}$ be an automorphism of $\mathcal{G}^{*}$ induced from the field automorphism $\varrho.$ Then $(\mathcal{G}^{*})^{\varrho^{*}}\cong{}^{2}{\rm E}_{6}(p^{b})_{ad},$ where $1\leq b<f,$ so $\ell=\ell_{18f}(p)$ does not divide $|(\mathcal{G}^{*})^{\varrho^{*}}|,$ and hence by [8, Lemma 2.5(ii)], $\chi_{s}$ is not $\varrho$-invariant. As $\chi_{s}$ is invariant under the diagonal automorphisms, $\chi_{s}$ is not $\sigma$-invariant and so some proper multiple of $\chi_{s}(1)$ is a degree of $H,$ which is impossible.
Subcase $H\cong\textrm{E}_{6}(q).$ Note that $\textrm{Out}(H)\cong\langle\gamma\rangle(\langle\varrho\rangle\times\langle\tau\rangle))$ is a group of order $2df,$ where $d=(3,q-1),$ and $\gamma$ is a diagonal automorphism of order $d,$ and $\varrho$ is a field automorphism of order $f$ and $\tau$ is a graph automorphism of order $2$ (see for example [10, Theorem 2.5.12]). Moreover $\tau$ inverts the cyclic group $\langle\gamma\rangle.$
Assume first that $G/C$ possesses only inner and diagonal automorphism. It follows that $G/C\cong{\rm E}_{6}(q)_{ad},$ $3\mid(q-1)$ and $\sigma$ is a diagonal automorphism of $H.$ Thus ${\mathrm{cd}}({\rm E}_{6}(q)_{ad})={\mathrm{cd}}(G/C)\subseteq{\mathrm{cd}}(G)={\mathrm{cd}}(H),$ which is impossible by Lemma 3.3$(x)$.
Assume next that $\sigma=\gamma\varrho,$ where $\varrho$ is a nontrivial field automorphism and $\gamma$ is either trivial or a diagonal automorphism. Let $\ell=\ell_{12f}(p)$ be a primitive prime divisor of $p^{12f}-1$ and let $s$ be a semisimple element of order $\ell$ in $K\cong{\rm F}_{4}(q)$ where $K\leq(H^{*})^{\prime}$ is chosen so that it is fixed under the involutory graph $\tau$ (see [10, Proposition 4.9.2]). Now $|C_{(\mathcal{G}^{*})^{F^{*}}}(s)|=\Phi_{12}\Phi_{3}$ and $\mathcal{G}^{F}$ possesses an irreducible character $\chi_{s}$ with degree
$|(\mathcal{G}^{*})^{F^{*}}:C_{(\mathcal{G}^{*})^{F^{*}}}(s)|_{p^{\prime}}=|{\rm E}_{6}(q)_{ad}|_{p^{\prime}}/\Phi_{12}\Phi_{3}$
which is trivial on the center as $s\in(H^{*})^{\prime}$ so $\chi_{s}\in{\mathrm{Irr}}(H).$ Let $\varrho^{*}$ be an automorphism of $\mathcal{G}^{*}$ induced from the field automorphism $\varrho.$ Then $(\mathcal{G}^{*})^{\varrho^{*}}\cong{\rm E}_{6}(p^{b})_{ad},$ where $1\leq b<f,$ so $\ell=\ell_{12f}(p)$ does not divide $|(\mathcal{G}^{*})^{\varrho^{*}}|$. By [8, Lemma 2.5(ii)], $\chi_{s}$ is not $\varrho$-invariant. As $\chi_{s}$ is invariant under the diagonal automorphisms, it is not $\varrho$-invariant and hence some proper multiple of $\chi_{s}(1)$ is a degree of $H,$ which is impossible.
@@
Now assume that $\sigma=\gamma\varrho\tau,$ where $\gamma$ is either trivial or a diagonal automorphism, $\varrho$ is a nontrivial field automorphism and $\tau$ is a graph automorphism. As the field and graph automorphisms commute by [10, Theorem 2.5.12], we can write $\sigma=\gamma\tau\varrho.$ Choose $s\in(H^{*})^{\prime}$ as in previous paragraph with $|s|=\ell_{12}(q).$ Since $s$ in $\tau$-invariant, [26, Corollary 2.5] yields that $\chi_{s}^{\tau}=\chi_{\tau(s)}=\chi_{s}.$ As $\chi_{s}$ is also $\gamma$-invariant, by applying [26, Corollary 2.5] again $\chi_{s}$ is $\gamma\tau$-invariant. Thus $\chi_{s}^{\sigma}=\chi_{s}^{\gamma\tau\varrho}=\chi_{s}^{\varrho}\neq\chi_{s}.$ Therefore $\chi_{s}$ is not $\sigma$-invariant, which leads to a contradiction.